OpenStudy (anonymous):

Solve the system of linear equations below. y = -2x + 19 y = x + 7

OpenStudy (welshfella):

First multiply the second equation by -2 then adding this to the first equation will give you an equation in y only Solve this for y

OpenStudy (welshfella):

sorry - a mistake there - multiply the second equation by 2 not -2

OpenStudy (welshfella):

y = -2x + 19 2y = 2x + 14

OpenStudy (anonymous):

Ty : )

OpenStudy (welshfella):

can you add these 2 equations

OpenStudy (welshfella):

yw

OpenStudy (radar):

Another option, note how both equations are = to y.........so they would be equal to each other. You could write: -2x + 19 = x + 7 Now simple algebra will let you solve for x. Then substitute the value of obtained for x and solve for y.

OpenStudy (triciaal):

The first method is the elimination/addition method and the second is substitution method. There are two other methods as well by graphing and by matrix. After you find the value for one variable use substitution in the expression for the other variable to get the other value. For whichever method you use it is a good practice to check your solution (x, y) in each of the original because it is very common to make arithmetic errors.

OpenStudy (ybarrap):